OpenStudy (yanni):

-2(y+1)=7y-5+3(2y+2)

OpenStudy (anonymous):

first distribute: -2y+-2=7y-5+6y+6 combine like terms -2y+2=13y+1 15y=1 y=1/15

OpenStudy (yanni):

its actually -1/5

OpenStudy (yanni):

but thx

OpenStudy (anonymous):

lol oops you're right sorry its been a while since i've done this